# SpikeLLM: Scaling up Spiking Neural Network to Large Language Models via Saliency-based Spiking

### v1

@@ Abstract
The recent advancements in large language models (LLMs) with billions of parameters have significantly boosted their performance across various real-world applications. However, the inference processes for these models require substantial energy and computational resources, presenting considerable deployment challenges. In contrast, human brains, which contain approximately 86 billion biological neurons, exhibit significantly greater energy efficiency compared to LLMs with a similar number of parameters. Inspired by this, we redesign 7~70 billion parameter LLMs using bio-plausible spiking mechanisms, emulating the efficient behavior of the human brain. We propose the first spiking large language model as recent LLMs termed SpikeLLM. Coupled with the proposed model, a novel spike-driven quantization framework named Optimal Brain Spiking is introduced to reduce the energy cost and accelerate inference speed via two essential approaches: first (second)-order differentiation-based salient channel detection, and per-channel salient outlier expansion with Generalized Integrate-and-Fire neurons. Our proposed spike-driven quantization can plug in main streams of quantization training methods. In the OmniQuant pipeline, SpikeLLM significantly reduces 25.51% WikiText2 perplexity and improves 3.08% average accuracy of 6 zero-shot datasets on a LLAMA2-7B 4A4W model. In the GPTQ pipeline, SpikeLLM realizes a sparse ternary quantization, which achieves additive in all linear layers. Compared with PB-LLM with similar operations, SpikeLLM also exceeds significantly. We will release our code on GitHub.
@@ 1 Introduction
Recent Artificial Neural Networks (ANNs) have shown scaling up Large Language Models (LLMs) [4, 49, 57, 24] can be one of the most potential techniques to access Artificial General Intelligence. However, despite these unprecedented and promising achievements, steering LLMs imposes a tremendous burden in terms of energy costs and computational requirements. For instance, running inference on the LLAMA-2 70B model requires three A100-80 GPUs, and each energy consumption is 400W. This creates a significant obstacle for generalizing LLMs to real-world applications, especially where limited battery capacity and memory size are critical, such as in mobile devices. To lower these barriers and broaden the applications of LLMs, we focus on energy-efficient artificial intelligence. Besides the traditional model compression algorithms [15, 18], solutions still largely remain open.
@@
Compared with ANN-based LLMs, human brain nervous systems achieve superior intelligence with much less energy consumption [17, 23] and a comparable number of neurons, approximately 86 billion. For several decades, the brain-inspired computing (BIC) field [36, 59] focuses on mimicking the biological nature of the human brain to develop more efficient and general AI algorithms [35] and physical platforms [45, 43, 40]. Among these, spiking neural networks (SNNs) [35, 17] are particularly notable for their biological plausibility and binary event-driven efficiency [55, 45]. SNNs can be viewed as ANNs that substitute spiking neuronal dynamics in each neuron. Despite their potential efficiency advantage, recent SNNs face two significant bottlenecks: (i) Firing Rate Encoding [6, 26, 9]. The firing rate encoding in existing SNNs fails to capture adequate semantic information efficiently. This limitation can hinder the generalization ability of LLMs, which heavily rely on acquired knowledge. (ii) Optimization. In direct training [38, 50], spiking dynamics are non-differentiable, and gradient estimation is inaccurate in backpropagation through time (BPTT). In ANN-SNN conversion [19, 20, 6, 26], it often requires much more inference steps to simulate ANNs, which is impractical for scaling up to LLMs. These challenges have kept SNNs relatively small (under 1 B parameters) and hinder their ability to scale to the complexity of the human brain.
This work aims to bridge the scale gap between recent SNNs and ANN-based LLMs or even part of the human brain nervous system with 86B neurons; and develop efficient LLMs as an alternative to traditional model compression. To achieve this goal, we start by comparing traditional quantization and spiking neuronal dynamics. Traditional quantization functions efficiently encode binary digits in one step but has built-in drawbacks for outliers or salient values in low-bit conditions [51, 27]. On the other hand, the most popular Integrate-and-Fire (IF) [30, 2] neurons can auto-regressively encode more semantic information by more spiking steps but are much more inefficient in the original rate encoding method. This work views the auto-regressive spiking neurons as generalized quantizers and as an alternative to quantizing activations, weights, and self-attentions (query and KV-caches) in LLMs. A Generalized Integrate-and-Fire (GIF) neuron is defined to auto-regressively quantize salient outliers and address the efficiency issue in IF. Based on GIF, spiking neurons own the capacity to express adequate semantic information which is essential to LLMs.
To drive spiking neurons, we propose an Optimal Brain Spiking framework to achieve salient-based quantization, which is a weight-activation generalization of the classic Optimal Brain Surgeon (OBS) [21]. Different from OBS, we detect salient channels in both activations and weights and allocate more spiking steps in GIF neurons for salient ones. After quantization, one salient channel is expanded as T channels by T spikeing steps, while most channels maintain one-step (or fewer-step) quantization. We divide and conquer activation and weight saliency [21] detection by different approximations of their Taylor expansion. In detail, the first-order gradient and second-order Hessian metrics are leveraged for activations and weights respectively. The Optimal Brain Spiking framework can cooperate with main streams of post-training quantization or calibration frameworks and a series of spike-driven quantized LLMs are built up termed SpikeLLM. For instance, for low-bit weight-activation quantization [47], we observe significant performance improvements in WikiText2 [37], C4 [42] and 6 zero-shot datasets. To further achieve fully additive linear layers like previous SNNs, we conduct weight-only quantization similar to GPTQ [16] pipeline. Compared with PB-LLM [46] in similar cost, SpikeLLM exceeds dramatically.
@@
  • We first scale up spiking neuronal dynamics to more than 10B parameters. Spiking large language models with 7~70 billion parameters are proposed, and could become the alternative to more bio-plausible and energy-efficient AI.
  • We propose a Generalized Integrate-and-Fire (GIF) neuron as a general alternative to traditional quantizers. The salient outliers are quantized auto-regressively and accurately.
  • We propose an Optimal Brain Spiking framework to drive the GIF quantizers. Per-channel saliency is efficiently detected with the first (second)-order differentiation-based metrics. Compared with state-of-the-art quantization frameworks, significant performance improvements are achieved with spiking neuronal dynamics.
@@ 2 Related Works
Brain-Inspired Computing. The Brain-Inspired Computing (BIC) [36, 59] field focuses on building up bio-plausible and general fundamental theories, algorithms, software [12], and hardware platforms [45, 43, 40] inspired by the structures and functions of biological nervous systems like the human brain. Spiking neural network (SNN) [43, 35] is one of the most popular BIC algorithms which embeds biological spiking neural dynamics in each single neuron [55, 45]. Promoted by the development of both deep learning and advanced biological neuron science, recent SNNs focus on the deep residual learning [13], self-attention [54, 61], normlization [60], as well as biological learning rules [39], structures [41] and energy efficiency [45]. In optimization, recent SNNs apply ANN-SNN conversion [19, 20, 6, 26] or directly training [38, 50] techniques. Most SNNs focus on the computation vision field; language-oriented SNNs are almost less than 1 billion parameters, for example, SpikeLM [52], SpikingBERT [1], SpikeBERT [33], and SpikeGPT [62].
Model Quantization. Model quantization aims at reducing the bit-width of weights or activations to accelerate network inference. Recent quantization includes Post-Training Quantization (PTQ) [51, 16], Quantization Aware Training (QAT) [32], and calibration training methods [47]. For small models, QAT methods achieve higher performance because of training from scratch, for example, LSQ [11], U2NQ [31]. For LLM quantization, PTQ methods including GPTQ [16], GPTQ-ada [22], SpQR [10], OWQ [25], AWQ [28], and PB-LLM [46] are weight-only quantization; SmoothQuant [51] and RPTQ [56] achieve weight-activation quantization. Besides, LLM-QAT [32], QA-LORA [53], and calibration based methods including Omniquant [47], QLLM [29], and AffineQuant [34] achieve higher performance.
@@ 3 Problem Formulation
To acquire quantized large language models, we start with recent low-bit quantization methods. Although traditional quantization is straight-through to encode weights and activations into binary digits, it is too simple to adapt to very low-bit conditions, resulting in a significant performance drop in LLMs, especially for activation quantization. Towards accurate and bio-plausible quantized LLMs, we replace direct quantization with spiking neuronal dynamics and pursue the first spiking large languages model inspired by advanced biological neuron science.
3.1 Low-Bit Quantization
This indicates the complexity and energy consumption of MAC operations are proportional to $M\times N$. Therefore, the number of operations in a quantized model can be measured using the arithmetic computation effort (ACE) metric [58], which is defined as $M\times N$ for a MAC operation between the M-bit weight and N-bit activation. Recent LLMs contain more than 10B ~100B parameters, making it an essential requirement to push not only weights but also activations to lower bit-width.
@@ 3.2 Limitations of Traditional Quantization
Traditional quantization is an ill-posed problem between bit-width and quantization error, especially for post-training LLMs. In low-bit cases, the performance drop is often caused by quantization errors of outliers and salient values. Previous work has shown that outliers in activations have magnitudes over 100 $\times$ larger than most values, and salient values in weight matrices significantly impact the results. To more precisely quantize these values, previous methods such as AWQ [28], SpQR [10], SmoothQuant [51], and Omniquant [47] have proposed corresponding mitigation strategies. However, these methods are constrained by the limitations of traditional quantization frameworks:
(i) weight-activation quantization makes it hard to avoid quantization errors in outliers. AWQ [28] uses per-channel quantization step sizes to smooth outlier channels; however, it can only be applied to weight-only quantization, which is often not enough for LLM compression.
(ii) Per-channel quantization is unfriendly to deployment. Since matrix multiplication is calculated per-token, per-channel quantization cannot be directly used to accelerate. As mitigation, SmoothQuant [51] and OmniQuant [47] rebalance the quantization difficulty of activations and weights; however, they do not directly eliminate the impact of outliers.
(iii) Mix-precision quantization is hardware unfriendly. SpQR [10] and PB-LLM [46] use mixed-precision quantization to avoid quantizing salient weights; however, mix-precision introduces difficulties in hardware deployment.
Based on these observations, there is an urgent requirement to explore weight-activation quantized LLMs and avoid the drawbacks of traditional quantization caused by outlier and salient values.
3.3 Spiking Neuronal Dynamics for Quantization
In this section, we first introduce the bio-inspired SNNs and then discuss the possibility of alternating traditional quantized LLMs with spiking LLMs. SNNs can be considered ANNs with the addition of biological spiking neuronal dynamics in each single neuron. Without loss of generality, the biological soma dynamics can be approximately modeled using the first- or higher-order differential equations. The Integrate-and-Fire (IF) neuron is a first-order approximation of soma dynamics, combining the advantages of bio-plausibility and efficiency, and can be represented as follows:
$\mathbf{v}(t)=\mathbf{v}(t-1)+\mathbf{x}^{(\ell-1)}(t)-\mathbf{s}^{(\ell)}(t)V_{th},$ (3)
$\mathbf{x}^{(\ell)}(t)=\mathbf{W}\mathbf{s}^{(\ell)\top}(t)V_{th}+\mathbf{W}\min{(\mathbf{x}^{(\ell-1)})}^{\top}+{b},$ (5)
where the IF neuron encodes a binary spike in each step $t$ for a duration of $T$. In Eq.3, the membrane potential $\mathbf{v}(t)$ accumulates current inputs $\mathbf{x}^{(\ell-1)}(t)$ to the last time step $\mathbf{v}(t-1)$ to simulate the charging process in soma. A subjection reset is applied to subsect the spiking values from the membrane potential $\mathbf{v}(t)$. In Eq.4, if $\mathbf{v}(t)$ exceeds a certain firing threshold $V_{th}$, the neuron is fired and encodes the spike $\mathbf{s}^{(\ell)}(t)$ as 1; otherwise, encodes as 0. Previous SNNs take the IF neuron as the activation quantizer. If we embed spiking neurons before a linear layer, this linear layer are derived as Eq.5 and the IF neurons perform asymmetric quantization. In Eq. 5, $\mathbf{W}\min{(\mathbf{x}^{(\ell-1)})}^{\top}+{b}$ formulates the bias of this linear layer.
Compared with the asymmetric uniform quantization in Eq.1 that encodes all binary digits at the same time, IF neurons auto-regressively encode per spike and can be viewed as a sequential expansion of the quantization function. In fact, IF neuron equals uniform quantization in some of the ANN-SNN conversion methods [5, 26]. In addition to previous ReLU activation-based conversion, we derive the equivalent forms of spiking linear towards asymmetric uniform quantization in Eq 5 and the quantized input of this layer can be represented as follows, where $\bar{\mathbf{s}}$ is the average of ${\mathbf{s}}(t)$:

### v2

@@ Abstract
Recent advancements in large language models (LLMs) with billions of parameters have improved performance in various applications, but their inference processes demand significant energy and computational resources. In contrast, the human brain, with approximately 86 billion neurons, is much more energy-efficient than LLMs with similar parameters. Inspired by this, we redesign 7$\sim$70 billion parameter LLMs using bio-plausible spiking mechanisms, emulating the efficient behavior of the human brain. We propose the first spiking large language model, SpikeLLM. Coupled with the proposed model, two essential approaches are proposed to improve spike training efficiency: Generalized Integrate-and-Fire (GIF) neurons to compress spike length from $T$ to $\frac{T}{L}\log_{2}L$ bits, and an Optimal Brain Spiking framework to divide outlier channels and allocate different $T$ for GIF neurons, which further compresses spike length to approximate $log_{2}T$ bits. The necessity of spike-driven LLM is proved by comparison with quantized LLMs with similar operations. In the OmniQuant pipeline, SpikeLLM reduces 11.01% WikiText2 perplexity and improves 2.55% accuracy of common scene reasoning on a LLAMA-7B W4A4 model. In the GPTQ pipeline, SpikeLLM achieves direct additive in linear layers, significantly exceeding PB-LLMs. Our code is publicly available at https://github.com/Xingrun-Xing2/SpikeLLM.
@@ 1 Introduction
Recent Artificial Neural Networks (ANNs) have shown scaling up Large Language Models (LLMs) (Brown et al., 2020; Touvron et al., 2023b; Zhang et al., 2022a; Le Scao et al., 2023) can be one of the most potential techniques to access Artificial General Intelligence. However, despite these unprecedented and promising achievements, steering LLMs imposes a tremendous burden in terms of energy costs and computational requirements. For instance, running inference on the LLAMA-2-70B model requires 2$\times$ A100-80 GPUs, and each energy consumption is 400W. This creates a significant obstacle for deploying LLMs to real-world applications, especially where limited battery capacity and memory size are critical, such as in mobile devices (Xing et al., 2025). To lower these barriers and broaden the applications of LLMs, we focus on energy-efficient artificial intelligence.
@@
Compared with ANN-based LLMs, human brain nervous systems achieve superior intelligence with much less energy consumption (Gerstner et al., 2014; Izhikevich, 2003) and a comparable number of neurons, approximately 86 billion. For several decades, the brain-inspired computing (BIC) field (Mehonic & Kenyon, 2022; Zhang et al., 2020) focuses on mimicking the biological nature of the human brain to develop more efficient and general AI algorithms (Maass, 1997) and physical platforms (Schuman et al., 2022; Roy et al., 2019; Pei et al., 2019). Among these, spiking neural networks (SNNs) (Maass, 1997; Gerstner et al., 2014) are particularly notable for their biological plausibility and binary event-driven efficiency (Yin et al., 2021; Schuman et al., 2022). Despite their potential efficiency advantage, recent SNNs face two significant bottlenecks: (i) Firing Rate Encoding (Bu et al., 2023; Li et al., 2021; Deng & Gu, 2021) in existing SNNs is inefficient in capturing adequate semantic information. As shown in Table 1 and Fig. 1 (a, b), to express $T$ quantization levels, a typical Integrate-and-Fire (IF) (Liu & Wang, 2001; Barbi et al., 2003) neuron requires $T$ time steps to encode full information, while quantization only requires $log_{2}T$ bit digits in one step. (ii) Inefficient Optimization. Direct training (Neftci et al., 2019; Wu et al., 2018) SNNs need gradient estimation in BackPropagation Through Time (BPTT) because of non-differentiable spiking dynamic; the ANN-SNN conversion (Han et al., 2020; Hao et al., 2023; Bu et al., 2023; Li et al., 2021) often requires much more inference steps to simulate ANNs, both of which are impractical for scaling up to LLMs. These challenges have kept SNNs relatively small (under 1 B parameters) and hinder their ability to scale to the complexity of the human brain.
This work aims to scale up SNNs to large language models or even part of the human brain nervous system with 86B neurons. Towards this goal, we propose both a micro spiking neuron design to improve spike efficiency and a macro saliency-based spiking design to drive spiking neurons. Currently, the primary approach to encode real-valued features to limited binary digits is model quantization. Given $T$ quantization levels, quantization functions can efficiently encode features into $log_{2}T$ binary digits in one step but encounter significant quantization errors for outliers or salient values in low-bit conditions (Xiao et al., 2023; Lin et al., 2023b). Inspired by both quantized ANNs and SNNs, we introduce a hybrid encoding method of quantization and spike firing rate encoding, termed Generalized Integrate-and-Fire (GIF) neurons. In each spiking step, we merge $L$ steps in IF neurons and encode as $log_{2}L$ bits as quantization; across different steps, we keep the recursive spiking dynamics. Compared with IF neurons, GIF compresses spike length from $T$ bits to $\frac{T}{L}log_{2}L$ bits, where $L$ is the quantization levels in each step. Compared with quantization, GIF neurons maintain recursive encoding advantages to accurately quantize salient channels with multisteps, which gives the potential to exceed quantization. To further determine channel-wise spiking steps in GIF neurons, as shown in Fig. 1 (d), a saliency-based spiking mechanism is proposed to divide and conquer salient channels and almost non-salient ones. And then, we allocates multistep spiking to encode salient channels and one-step spiking for others, which further compresses to approximate $log_{2}T$ bits over all channels. Unlike mix-precision quantization in Fig. 1 (c), saliency-based spiking expands salient channels with GIF neurons and equals to single-precision quantization in Fig. 1 (d).
Figure 1: Different encoding methods. In (a, b), the activation has N channels; each value has T quantization levels. Given salient channels (in blue), mix-precision methods (c) are deployment unfriendly. In spike-driven methods (d), we expand salient channels by spiking dynamics to realize single precision quantization, where T’ is spiking steps in salient channels.
Table 1: Comparason of encoding efficiency. L indicates quantization levels in each step, where $L\in[1,T]$. The accuracy and perplexity are evaluated in the common scene reasoning and WikiText with the LLAMA-7B, W4A4 setting (Appendix A.2).
method Steps Bits/Step Bits Quant-Levels Acc. $\uparrow$ PPL $\downarrow$
IF-SNN $T$ 1 $T$ $T$
GIF-SNN $\frac{T}{L}$ $log_{2}L$ $\frac{T}{L}log_{2}L$ $T$
Quant-ANN 1 $log_{2}T$ $log_{2}T$ $T$ 49.93 11.85
SpikeLLM $\approx 1$ $log_{2}T$ $\approx log_{2}T$ $\approx T$ 52.48 9.98
To detect salient channels, we propose a framework named Optimal Brain Spiking (OBSpiking), which is a weight-activation generalization of the classic Optimal Brain Surgeon (OBS) (Hassibi & Stork, 1992) for model pruning. The key concept of our method is distinguished approximations of weight and activation Taylor expansion to calculate their saliency. In detail, the first-order gradient and second-order Hessian metrics are leveraged for activations and weights respectively. For every matrix multiplication in LLMs, GIF neurons can be viewed as generalized quantizers to quantize salient and other channels based on the saliency rank generated by the OBSpiking framework.
To evaluate the necessity of the Spiking LLMs, we integrate the proposed spiking mechanism with the most classic LLM quantization pipelines including the Omniquant (Shao et al., 2023) and GPTQ, named SpikeLLM. For weight-activation quantization (Shao et al., 2023), we observe significant performance improvements in generation and common scene reasoning. For weight-only quantization, GIF neurons further quantize ternary weights with the GPTQ (Frantar et al., 2022) pipeline, achieving direct addition networks. SpikeLLM exceeds PB-LLM (Shang et al., 2023) dramatically.
@@
  • We first scale up spiking neuronal dynamics to 7$\sim$70 billion parameters, promoting SNNs to the era of LLMs. The necessity of introducing a spiking mechanism to LLMs is proved by the comparison with quantization methods.
  • We propose a Generalized Integrate-and-Fire neuron and Optimal Brain Spiking framework as a general alternative to traditional quantizers. The efficiency issue of firing rate encoding in SNNs is addressed by spike length compression.
  • The proposed GIF neuron and OBSpiking can be implemented by both general matrix multiplication (GEMM) and binary event-driven operations, providing the first practical design of SNNs with more than tens of billion parameters.
@@ 2 Related Works
Brain-Inspired Computing. The Brain-Inspired Computing (BIC) (Mehonic & Kenyon, 2022; Zhang et al., 2020) field focuses on building up bio-plausible and general fundamental theories, algorithms, software (Fang et al., 2023), and hardware platforms (Schuman et al., 2022; Roy et al., 2019; Pei et al., 2019) inspired by the structures and functions of biological nervous systems like the human brain. Spiking neural network (SNN) (Roy et al., 2019; Maass, 1997) is one of the most popular BIC algorithms which embeds biological spiking neural dynamics in each single neuron (Yin et al., 2021; Schuman et al., 2022). Promoted by the development of both deep learning and advanced biological neuron science, recent SNNs focus on the deep residual learning (Fang et al., 2021), self-attention (Yao et al., 2023; Zhou et al., 2023), normlization (Zheng et al., 2021), as well as biological learning rules (Payeur et al., 2021), structures (Pham et al., 2021) and energy efficiency (Schuman et al., 2022). In optimization, recent SNNs apply ANN-SNN conversion (Han et al., 2020; Hao et al., 2023; Bu et al., 2023; Li et al., 2021) or direct training (Neftci et al., 2019; Wu et al., 2018) techniques. Most SNNs focus on the computation vision field; language-oriented SNNs are almost less than 1 billion parameters, for example, SpikeLM (Xing et al., 2024b), SpikingBERT (Bal & Sengupta, 2024; Lv et al., 2023; Su et al., 2024), and SpikeGPT (Zhu et al., 2023). How to scale up bio-inspired spiking mechanisms to billions of parameters has become a valuable issue.
Model Quantization. Model quantization aims at reducing the bit-width of weights or activations to accelerate network inference exemplified by the recent quantization works including Post-Training Quantization (PTQ) (Xiao et al., 2023; Frantar et al., 2022), Quantization Aware Training (QAT) (Liu et al., 2023b), and calibration training methods (Shao et al., 2023). For small neural networks, QAT methods (Esser et al., 2019; Liu et al., 2022; Xing et al., 2024a; 2022b; 2022a), benefit from the specifically designed training protocol, achieving outstanding performance. However, these methods are not practical in the LLM setting. In contrast, PTQ methods including GPTQ (Frantar et al., 2022), GPTQ-ada (Heo et al., 2023), SpQR (Dettmers et al., 2023), OWQ (Lee et al., 2023), AWQ (Lin et al., 2023a), and PB-LLM (Shang et al., 2023) tailored for LLM are weight-only quantization; SmoothQuant (Xiao et al., 2023) and RPTQ (Yuan et al., 2023) achieve weight-activation quantization. Besides, LLM-QAT (Liu et al., 2023b), QA-LORA (Xu et al., 2023), and calibration based methods including Omniquant (Shao et al., 2023), QLLM (Liu et al., 2023a), AffineQuant (Ma et al., 2024), and QuaRot (Ashkboos et al., 2024) achieve higher performance. However, current quantized LLMs encounter significant quantization errors in outlier or salient channels as discussed in Section 3.2. Althrough more recent rotation quantization (Ashkboos et al., 2024) achieves outlier-free, there needs additional matrix multiplications that can not be absorbed.
@@ 3 Problem Formulation
In this section, we first introduce bio-inspired Spiking Neural Networks (SNNs) and explore the potential of replacing traditional quantized Large Language Models (LLMs) with spiking LLMs. Additionally, we address the challenges of outlier and salient value quantization, which motivate the introduction of the spiking mechanism.
3.1 Spiking Neuronal Dynamics
SNNs can be considered ANNs by adding biological spiking neuronal dynamics in each neuron. Without loss of generality, the biological soma dynamics are approximately modeled by the first- or higher-order differential equations. The IF neuron is a first-order approximation of soma dynamics, combining the advantages of bio-plausibility and efficiency, which can be represented by:
$\mathbf{v}(t)=\mathbf{v}(t-1)+[\mathbf{x}^{(\ell-1)}(t)-\min{(\mathbf{x}^{(\ell-1)})}]-\mathbf{s}^{(\ell)}(t-1)V_{th},$ (1)
$\mathbf{x}^{(\ell-1)}(t)=\mathbf{s}^{(\ell)}(t)V_{th}+\min{(\mathbf{x}^{(\ell-1)})},$ (3)
where the IF neuron encodes a binary spike in each step $t$ for a duration of $T$. In Eq.1, the membrane potential $\mathbf{v}(t)$ accumulates current input $\mathbf{x}^{(\ell-1)}(t)$ to the last time step $\mathbf{v}(t-1)$ to simulate the charging process in soma. A subjection reset is applied to subtract the spiking values from the membrane potential $\mathbf{v}(t)$. In Eq.2, if $\mathbf{v}(t)$ exceeds a certain firing threshold $V_{th}$, the neuron is fired and encodes the spike $\mathbf{s}^{(\ell)}(t)$ as 1; otherwise, encodes as 0. Thus, previous SNNs take the IF neuron as an activation quantizer, which encodes full-precision activation as 1-bit output per spiking step in Eq. 3. Different from SNNs with ReLU activations, to encode the negative values in transformers, we encode $\mathbf{x}^{(\ell-1)}(t)-\min{(\mathbf{x}^{(\ell-1)})}$ by spike firing rate, where $\min{(\mathbf{x}^{(\ell-1)})}$ is the min value of this activation token, and can be viewed as the zero-point (Shao et al., 2023) of the quantizer.
As shown in Table 1, compared with the asymmetric uniform quantization (Appendix A.1) that encodes $T$ levels via $log_{2}T$ bits, IF neurons recursively encode spikes via total $T$ bits, because firing rate encoding directly makes an average of spiking steps, which missing the numerical carry. In ANN-SNN conversion (Bu et al., 2021; Li et al., 2021), IF neuron equals uniform quantization, where SNNs expand $T$ steps of their $log_{2}T$ bit quantized ANN counterpart. By summation over spiking steps, the IF neuron encodes the input into $log_{2}T$ bit integer values, where $\bar{\mathbf{s}}$ is the spike firing rate; Clip and Round indicate the min-max clipping and floor functions:
@@ 3.2 Limitations of Traditional Quantization
Traditional quantization is an ill-posed problem between bit-width and quantization error, especially for post-training LLMs. The performance drop is often caused by quantization errors of outliers (Xiao et al., 2023) and salient values (Dettmers et al., 2023). In details, the traditional LLM quantization faces the following challenges:
(i) weight-activation quantization makes it hard to avoid quantization errors in outliers. AWQ (Lin et al., 2023a) uses per-channel quantization step sizes to smooth outlier channels; however, it can only be applied to weight-only quantization, which is often insufficient for LLM compression.
(ii) Per-channel quantization is unfriendly to deployment. Since matrix multiplication is calculated per-token, per-channel quantization cannot be directly used to accelerate. As mitigation, SmoothQuant (Xiao et al., 2023) and OmniQuant (Shao et al., 2023) rebalance the quantization difficulty of activations and weights; however, they do not directly eliminate the impact of outliers.
(iii) Mix-precision quantization is hardware unfriendly. SpQR (Dettmers et al., 2023) and PB-LLM (Shang et al., 2023) use mixed-precision quantization to avoid quantizing salient weights; however, mix-precision introduces difficulties in hardware deployment.
Additionally, recent QuaRot (Ashkboos et al., 2024) achieves outlier free quantization by rotation. The Hadamard transformation in Quarot maps activations to roughly ball-shape Gaussian distribution, where more accurate spike encoding for larger values still benefits performance in our experiments. Moreover, compared with traditional quantization, QuaRot adds 4 Hadamard matrixes in each bloch, and their multiplications could not be absorbed. Based on these challenges, there is a requirement to explore spike encoding methods which helps accurate quantization.